OpenStudy (anonymous):

SOMEBODY HELP!!!! QUICK!! Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

@jim_thompson5910 Can you please help me?

jimthompson5910 (jim_thompson5910):

Divide both sides by -5 to go from −5|x + 1| = 10 to |x+1| = -2

jimthompson5910 (jim_thompson5910):

what issue do you see with |x+1| = -2

OpenStudy (anonymous):

that it is equal to a negative?

jimthompson5910 (jim_thompson5910):

why is that an issue?

OpenStudy (anonymous):

bcuz i think it needs to be a positive and then i can do another problem that is a negative

jimthompson5910 (jim_thompson5910):

|x+1| is never negative because absolute value represents distance negative distance makes no sense at all

jimthompson5910 (jim_thompson5910):

so |x+1| = -2 has no solutions

OpenStudy (anonymous):

ooooooooooooo ok

OpenStudy (anonymous):

THANKS!!! I GET IT

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@jim_thompson5910 I like the way you teach, you sir are a most have on this forum

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking and thanks magepker728